# Strong Confinement of a Nanoparticle in a Needle Paul Trap: Towards Matter-Wave Interferometry with Nanodiamonds

### v1

@@ Abstract
Quantum mechanics (QM) and General relativity (GR), also known as the theory of gravity, are the two pillars of modern physics. A matter-wave interferometer with a massive particle, can test numerous fundamental ideas, including the spatial superposition principle - a foundational concept in QM - in completely new regimes, as well as the interface between QM and GR, e.g., testing the quantization of gravity. Consequently, there exists an intensive effort to realize such an interferometer. While several paths are being pursued, we focus on utilizing nanodiamonds as our particle, and a spin embedded in the nanodiamond together with Stern-Gerlach forces, to achieve a closed loop in space-time. There is a growing community of groups pursuing this path [1]. We are posting this technical note (as part of a series of seven such notes), to highlight our plans and solutions concerning various challenges in this ambitious endeavor, hoping this will support this growing community. In this work, we achieve strong confinement of a levitated particle, which is crucial for angular confinement, precise positioning, and perhaps also advantageous for deep cooling. We designed a needle Paul trap with a controllable distance between the electrodes, giving rise to a strong electric gradient. By combining it with an effective charging method - electrospray - we reach a trap frequency of up to 40 kHz, which is more than twice the state of the art. We believe that the designed trap could become a significant tool in the hands of the community working towards massive matter-wave interferometry. We would be happy to make more details available upon request.
@@ 4 Experiment
Figure 2: The experimental setup. A. The scheme of the setup. The needles are fixed inside an octagon vacuum chamber on two linear motorized piezo stages (Newport CONEX-AG-LS25-27P), connected to the controllers via vacuum feedthroughs. One of the stages controls the distance between the needles, and another controls the position of both needles. The trap is powered by a signal generator and a voltage amplifier (TREK model 2210). An infrared laser beam ($\lambda=1550\rm\,nm$) of total power $P\simeq 50\,\rm mW$ is focused into the trap with lens 1 (NA = 0.5). The scattered and unscattered light are collimated by lens 2 (NA = 0.39) and then split by a D-shaped mirror into two inputs of a balanced photodiode. The signal is analyzed with an NI PCI-5922 board serving as a spectrum analyzer. A green laser is used to illuminate the particle and detect it with a usual visible-range camera. B. Photo of the trap inside the octagon chamber. C. Electrospray system. The electrospray system (MolecularSpray UHV4i) is assembled vertically, so that the needle emitting the diamond solution is on top. The high voltage is applied to the solution, so that the charged particles fly from the needle through the grounded entrance capillary (internal diameter 0.25 mm) inside the vacuum. Entrance capillary leads to the ”Stage 1” which is separated from the octagon vacuum chamber with an additional differential-pumping aperture and is connected to a separate rotary pump.
@@ 5 Results and discussion
Figure 3: Particle’s motion spectrum at different distances between the needles. A. Power spectral density (PSD) of the split-detection signal of a diamond nanoparticle at three different distances between the needles $d$. The particle (Adamas Nanotechnologies, 40 nm nominal diameter) was trapped at a fixed RF voltage and frequency $V_{0}=326\,\rm V_{pp},$ $\Omega_{rf}=2\pi\times 114\,\rm kHz$ and at a fixed pressure of $0.2\rm\,Torr$. The distance between the needles was scanned using the piezo stage 1. The secular-motion peaks are clearly visible; all the other sharper peaks correspond to noise independent of the trap’s parameters. B. Maximum trap frequency as a function of the distance between the needles. The model takes into account higher-order corrections beyond the pseudopotential approximation according to [67], because $q_{z}$ reaches a value of 0.75 at the smallest distances. The dependence $\eta(d)$ is also taken into account. Once the mass is determined from the PSD peak width, the model has two free parameters: the charge of the particle estimated by the fit as $Q=4.85e$, and DC-shift of the applied voltage, estimated to be $U_{0}=9\,\rm V$ (the latter coincides with the measured value). See Appendix for the model’s details. We observed jumps of the trap frequency between two values separated by approximately 1 kHz, likely corresponding to two stable equilibrium positions, but further study is needed. The graph shows the highest-frequency mode.
Figure 3B shows the trap frequency of a diamond particle (Adamas Nanotechnologies, 40 nm nominal diameter) as a function of the distance between the needles $d$. As can be seen, we reached a trap frequency of more than 30 kHz (another particle reached 40 kHz), which is, to the best of our knowledge, at least twice as large as the state-of-the-art value for nanoparticles levitated in Paul traps [49]. The diameter of the particle, estimated from the PSD peak width [43], equals approximately 52.5 nm, which is within the manufacturer’s uncertainty. These values allow us to estimate the particle’s minimal uncertainty in position to be $\sigma_{z}\approx\sqrt{\frac{k_{B}T}{m\omega_{trap}^{2}}}\approx\rm 490\,nm$ along the axis of the strongest confinement.
@@ Acknowledgments
We thank the BGU Atom-Chip Group support team, especially Menachem Givon, Zina Binstock, Dmitrii Kapusta and Michael Averbukh for their support in building and maintaining the experiment. We also thank Or Dobkowski and Omer Feldman from our group, and Marko Toroš from the University of Ljubljana, for useful discussions. Funding: This work was funded by the Gordon and Betty Moore Foundation (doi.org/10.37807/GBMF11936), and Simons Foundation (MP-TMPS-00005477).
@@
This approximation gives an error of below 2% for our parameters. Eq. 10 was used to fit the experimental data in Fig. 3 with two fitting parameters: the charge of the particle and DC offset, returning the values $Q=4.85e$ and $U_{0}=9\rm\,V$ (the latter coincides with the separately measured value).

### v2

# Strong Confinement of a Nanodiamond in a Needle Paul Trap: Towards Matter-Wave Interferometry with Massive Objects

@@ Abstract
Quantum mechanics (QM) and General relativity (GR), also known as the theory of gravity, are the two pillars of modern physics. A matter-wave interferometer with a massive particle, can test numerous fundamental ideas, including the spatial superposition principle - a foundational concept in QM - in completely new regimes, as well as the interface between QM and GR, e.g., testing the quantization of gravity. Consequently, there exists an intensive effort to realize such an interferometer. While several paths are being pursued, we focus on utilizing nanodiamonds as our particle, and a spin embedded in the nanodiamond together with Stern-Gerlach forces, to achieve a closed loop in space-time. There is a growing community of groups pursuing this path [1]. We are posting this technical note (as part of a series of seven such notes), to highlight our plans and solutions concerning various challenges in this ambitious endeavor, hoping this will support this growing community. In this work, we achieve strong confinement of a levitated particle, which is crucial for angular confinement, precise positioning, and also advantageous for deep cooling. We designed a needle Paul trap with a controllable distance between the electrodes, giving rise to a strong electric gradient. By combining it with an effective charging method - electrospray - we reach a trap frequency of up to 40 kHz, which is more than twice the state of the art. We believe that the designed trap could become a significant tool in the hands of the community working towards massive matter-wave interferometry. We would be happy to make more details available upon request.
@@ 4 Experiment
Figure 2: The experimental setup. A. The scheme of the setup. The needles are fixed inside an octagon vacuum chamber on two linear motorized piezo stages (Newport CONEX-AG-LS25-27P), connected to the controllers via vacuum feedthroughs. One of the stages controls the distance between the needles, and another controls the position of both needles. The trap is powered by a signal generator and a voltage amplifier (TREK model 2210). An infrared laser beam ($\lambda=1550\rm\,nm$) of total power $P\simeq 50\,\rm mW$ is focused into the trap with lens 1 (NA = 0.5). The scattered and unscattered light are collimated by lens 2 (NA = 0.39) and then split by a D-shaped mirror into two inputs of a balanced photodiode. The signal is analyzed with an NI PCI-5922 board serving as a spectrum analyzer. A green laser is used to illuminate the particle. The particle is detected with a visible-range or infrared (IR) camera. B. Photo of the trap inside the octagon chamber. C. Electrospray system. The electrospray system (MolecularSpray UHV4i) is assembled vertically, so that the needle emitting the diamond solution is on top. The high voltage is applied to the solution, so that the charged particles fly from the needle through the grounded entrance capillary (internal diameter 0.25 mm) inside the vacuum. Entrance capillary leads to the ”Stage 1” which is separated from the octagon vacuum chamber with an additional differential-pumping aperture and is connected to a separate rotary pump.
@@ 5 Results and discussion
Figure 3: Particle’s motion spectrum at different distances between the needles. A. Power spectral density (PSD) of the split-detection signal of a diamond nanoparticle at three different distances between the needles $d$. The particle (Adamas Nanotechnologies, 40 nm nominal diameter) was trapped at a fixed RF voltage and frequency $V_{0}=326\,\rm V_{pp},$ $\Omega_{rf}=2\pi\times 114\,\rm kHz$ and at a fixed pressure of $0.2\rm\,Torr$. The distance between the needles was scanned using the piezo stage 1. The secular-motion peaks are clearly visible; all the other sharper peaks correspond to noise independent of the trap’s parameters. B. Maximum trap frequency as a function of the distance between the needles. The model takes into account higher-order corrections beyond the pseudopotential approximation according to [67], because $q_{z}$ reaches a value of 0.75 at the smallest distances. The dependence $\eta(d)$ is also taken into account. The model has two free parameters: the charge-to-mass ratio of the particle estimated by the fit as $Q/m=2.92\,\rm C/kg$, and DC-shift of the applied voltage, estimated to be $U_{0}=9\,\rm V$ (the latter coincides with the independently measured value). See Appendix for the model’s details. We observed jumps of the trap frequency between two values separated by approximately 1 kHz, likely corresponding to two stable equilibrium positions, but further study is needed. The graph shows the highest-frequency mode.
Figure 3B shows the trap frequency of a diamond particle (Adamas Nanotechnologies, 40 nm nominal diameter) as a function of the distance between the needles $d$. As can be seen, we reached a trap frequency of more than 30 kHz (another particle reached 40 kHz), which is, to the best of our knowledge, at least twice as large as the state-of-the-art value for nanoparticles levitated in Paul traps [49]. The diameter of the particle, estimated from the PSD peak width [43], equals $63.6\pm 3.8$ nm, which is within the manufacturer’s uncertainty. These values allow us to estimate the particle’s minimal uncertainty in position to be $\sigma_{z}\approx\sqrt{\frac{k_{B}T}{m\omega_{trap}^{2}}}\approx\rm 370\,nm$ along the axis of the strongest confinement.
@@ Acknowledgments
We thank the BGU Atom-Chip Group support team, especially Menachem Givon, Zina Binstock, Dmitrii Kapusta and Michael Averbukh for their support in building and maintaining the experiment. We also thank Or Dobkowski and Omer Feldman from our group, and Marko Toroš from the University of Ljubljana, for useful discussions. Funding: This work has been supported by the ”Table-top experiments for fundamental physics” program, sponsored by the Gordon and Betty Moore Foundation, Simons Foundation, Alfred P. Sloan Foundation, and John Templeton Foundation.
@@
This approximation gives an error of below 2% for our parameters. Eq. 10 and eq. 9 were used to fit the experimental data in Fig. 3 with two fitting parameters: the charge-to-mass ratio of the particle and DC offset, returning the values $Q/m=2.92\,\rm C/kg$ and $U_{0}=9\rm\,V$ (the latter coincides with the separately measured value).